a. From the information given here determine the 95% confidence interval estimate of the population mean.

x = 100  = 20 n = 25

b. Repeat part (a) with x = 200.

c. Repeat part (a) with x = 500.

d. Describe what happens to the width of the confidence interval estimate when the sample mean increases.


1
Expert's answer
2020-12-14T13:08:10-0500

"\bar x=100"

"\sigma =20"

"n =25"


a) Confidence interval is:


"\bar x \pm Z_{\frac{\alpha}{2}}\frac{\sigma}{\sqrt n}"

"100\pm Z_{\frac{1-0.95}{2}}\frac{20}{\sqrt{25}}=100\pm Z_{0.025}\cdot 4=100\pm1.96 \cdot4=100\pm 7.84"

So given the 95% confidence level the interval is (92.16, 107.84)

b) "\bar x=200": "200\pm 7.84"

The 95% confidence interval is (192.16, 207.84)

c) "\bar x=500": "500\pm 7.84"

The 95% confidence interval is (492.16, 507.84)

d) As we can see, increasing the sample mean does not change the width of the confidence interval; it only changes where the interval is centered.